# Non-linear conductances of Galton–Watson trees and application to the (near) critical random cluster model

### v1

@@ 5.1 Convergence in $L^{1}$
Let us write $f(x)=x-g(x)$, which is a non-negative function (by concavity of $g$), which verifies $f(x)\leqslant c\min(x^{q},x)$ thanks to (1.5). Then, we can rewrite the above iteration as
@@
which goes to 00 as $k\to\infty$ since $\theta_{q}<q$.
@@
where we have used Lemma 4.3. We now show that the upper bound goes to 00 as $k\to\infty$.
@@
which goes to 00 if $k\to\infty$ sufficiently slowly.
@@
The case $mR_{n}=1$ is trivial since then we have $a_{k}=k$ and $a_{n}=n$. A similar result holds when $|mR_{n}-1|\leqslant n^{-1}$, since we also have $a_{k}\asymp k$ for all $k\leqslant n$ in that case.
@@
For the last inequality, we have used that $f(x)\leqslant c\,\min(x^{q},x)$, so that for $|v|=j$ we have $\mathbf{E}[f(B_{n}(v))]\leqslant c\mathbf{E}[B_{n}(v)^{q}]\leqslant c^{\prime}(b_{n-j})^{q}$, thanks to Proposition 4.2; recall also that $s=q-1$.
@@
$\begin{split}\mathbf{E}\big{[}|\hat{B}_{n}-W|\big{]}&\leqslant\mathbf{E}\big{[}|W_{k}-W|\big{]}+cm^{-k(1-\frac{1}{q}\theta_{q})}+c\,\frac{a_{k}}{a_{n}}\\ &\leqslant c^{\prime}m^{-k(1-\frac{1}{q}\theta_{q})}+c\,\frac{a_{k}}{a_{n}}\,,\end{split}$ (5.6)
For the last inequality, we have used that $\mathbf{E}[|W-W_{k}|]\leqslant\|W-W_{k}\|_{q}\leqslant cm^{-k(1-\frac{1}{q}\theta_{q})}$, see [24, Prop. 1.3].
Note that the assumption that $\limsup_{n\to\infty}mR_{n}\leqslant 1$ ensures that $k_{n}\to\infty$ and that $a_{n}\to+\infty$, see (1.10). Hence, we can choose $k=\hat{k}_{n}\leqslant k_{n}$ going to +$\infty$+ ∞ sufficiently slowly so that the upper bound in (5.6) goes to zero. This concludes the proof that $(\hat{B}_{n})_{n\geqslant 0}$ converges in $L^{1}$ to $W$. ∎
@@ 5.2 Convergence in $L^{q}$
First, (5.3) and (5.4) show that $(I_{\hat{k}_{n},n})^{q}$ and $(J_{\hat{k}_{n},n})^{q}$ converge to 00 in $L^{1}$; in particular they are uniformly integrable. Second, since $\mu$ admits a finite moment of order $q$, we have that $(W_{k})_{k\geqslant 0}$ converges in $L^{q}$ to $W$, so in particular $(W_{k}^{q})_{k\geqslant 0}$ is uniformly integrable. This concludes the proof that $(\hat{B}_{n}^{q})_{n\geqslant 0}$ is uniformly integrable, hence that $(\hat{B}_{n})_{n\geqslant 0}$ converges in $L^{q}$ to $W$. ∎
@@ 5.3 Almost sure convergence
Note that, setting $f(x)=x-g(x)$, we have the relation $b_{n}=mR_{n}\big{(}b_{n-1}-\mathbf{E}[f(B_{n}(v))]\big{)}$ for $|v|=1$, which can be rewritten as
@@
Now, observe that since $f(x)\sim\kappa_{g}x^{q}$ as $x\downarrow 0$ and that $f(x)\leqslant cx^{q}$, we get by dominated convergence that $\mathbf{E}[f(B_{n}(v))]\sim\kappa_{g}\mathbf{E}[B_{n}(v)^{q}]\sim\kappa_{g}\mathbf{E}[W^{p}](b_{n-1})^{p}$, with $b_{n-1}\downarrow 0$ under the assumption of Proposition 1.9.
In particular, this gives that $c_{n-1}\sim\kappa_{g}\mathbf{E}[W^{p}](b_{n-1})^{s}$, with $s=q-1$. Setting $x_{n}=b_{n}^{-s}$, we get that
@@
with $c^{\prime}_{n-1}=(1-c_{n-1})^{-s}-1\sim s\kappa_{g}(b_{n-1})^{s}$. Iterating this relation, we get that
@@
with $x_{i}c^{\prime}_{i}\to s\kappa_{g}\mathbf{E}[W^{q}]$ as $i\to\infty$. Hence, we can choose $\ell_{n}$ going to +$\infty$+ ∞ arbitrarily slowly and write that
$\begin{split}x_{n}&=(1+o(1))s\kappa_{g}\mathbf{E}[W^{p}]\sum_{k=1}^{n-\ell_{n}}(mR_{n})^{-ks}+(mR_{n})^{-ns}\sum_{i=0}^{\ell_{n}}(mR_{n})^{is}x_{i}c^{\prime}_{i}\\ &=(1+o(1))s\kappa_{g}\mathbf{E}[W^{p}]a_{n-\ell_{n}}+(mR_{n})^{-ns}\sum_{i=0}^{\ell_{n}}(mR_{n})^{is}x_{i}c^{\prime}_{i}\,.\end{split}$
To conclude, let us notice that, in the case where $\lim_{n\to\infty}mR_{n}=\vartheta=1$, then we can choose $\ell_{n}\to\infty$ so that $a_{n-\ell_{n}}\sim a_{n}$. We also find in that case that the second term is negligible compared to $a_{n}$ provided that $\ell_{n}$ grows sufficiently slowly — this is clear if $mR_{n}\geqslant 1-\frac{c}{n}$ and follows from the fact that $a_{n}\sim(1-(mR_{n})^{s})^{-1}(mR_{n})^{-(n+1)s}$ if $n(mR_{n}-1)\to-\infty$. This shows that $x_{n}\sim s\kappa_{g}\mathbf{E}[W^{p}]a_{n}$ as $n\to\infty$ when $\vartheta=1$.
@@ 6 Estimates on moments of $C_{n}$ in the case $\mathbf{E}[Z^{q}]=+\infty$
In this section, we prove Theorem 1.10. Since we can bound $B_{n}$ with $C_{n}$, see Proposition 1.4, we focus on estimates on $C_{n}$, $\mathbf{E}[C_{n}]$. Let us define, as in Section 4.1.1, $\phi(u)=R_{u}C_{n}(u)$ and recall that we have the following recursion (4.1):
@@ 6.1 Upper bound on $\mathbf{E}[C_{n}]$
Let us assume that (1.12) holds and obtain the upper bound in Theorem 1.10. As in Section 4.2, let $\varphi_{k}=\mathbf{E}[\phi(u)]$ for $|u|=k$, so in particular $\varphi_{n}=1$ and we need to estimate $\varphi_{0}$. Using the recursion (6.1) we can write, again with the notation $f(x)=x-g(x)$,
$\begin{split}\mathbf{E}[g(\phi(u))]&=mR_{n}\mathbf{E}[g(\phi(v))]-\mathbf{E}\Big{[}f\Big{(}R_{n}\sum_{v\leftarrow u}g\big{(}\phi(v)\big{)}\Big{)}\Big{]}\\ &=mR_{n}\mathbf{E}[g(\phi(v))]-\mathbf{E}\Big{[}f\Big{(}R_{n}Z_{u}\mathbf{E}\big{[}g\big{(}\phi(v)\big{)}\big{]}\Big{)}\Big{]}\\ &\qquad\qquad\qquad\qquad\qquad-\mathbf{E}\bigg{[}f\Big{(}R_{n}\sum_{v\leftarrow u}g\big{(}\phi(v)\big{)}\Big{)}-f\Big{(}R_{n}Z_{u}\mathbf{E}\big{[}g\big{(}\phi(v)\big{)}\big{]}\Big{)}\bigg{]}\,,\end{split}$ (6.2)
@@
Notice that the recursion (6.4) admits a non-zero fixed point $u_{*}$ if $mR_{n}>1$, which verifies $h(u_{*})=\frac{mR_{n}-1}{mR_{n}}$, but that if $mR_{n}\leqslant 1$ the only fixed point is $u_{*}=0$.
• Let us start with the case where $mR_{n}\in[1+\frac{1}{n},2]$, and let $\delta_{n}:=\frac{mR_{n}-1}{mR_{n}}$. We let $\upsilon_{n}$ be the fixed point of the equation (6.4), i.e. such that $h(\upsilon_{n})=\delta_{n}$, and notice that it verifies $\upsilon_{n}=h^{-1}(\delta_{n})\leqslant c\gamma_{n}$, where $\gamma_{n}$ is defined in (1.14). Let assume that $\upsilon_{n}\leqslant u_{0}$, otherwise we have $u_{k}\leqslant\upsilon_{n}$ for all $k$ and in particular $u_{n}\leqslant\upsilon_{n}\leqslant c\gamma_{n}$.
@@
Now, applying this inequality with $k=k_{n}-1$ and recalling the definition of $k_{n}$, we get that $k_{n}-1\leqslant c^{\prime\prime}C^{-1}/\delta_{n}$. Since $\delta_{n}=(mR_{n}-1)/R_{n}\geqslant 1/2n$ for $mR_{n}\in[1+\frac{1}{n},2]$, we get that $k_{n}-1\leqslant 2c^{\prime\prime}C^{-1}n$, which is smaller than $n$ provided that $C$ had been fixed large enough.
@@ 6.2 Lower bound on $C_{n}$, $\mathbf{E}[C_{n}]$
To obtain a lower bound on $C_{n}$, $\mathbf{E}[C_{n}]$, let us assume that (1.13) holds. We let $(t_{n})_{n\geqslant 1}$ be a truncation sequence (to be optimized later on), and as in 3.2, we consider a truncated branching process $\tilde{T}$ with offspring distribution $\tilde{Z}:=(Z\wedge t_{n})\sim\tilde{\mu}$. Let also $\tilde{C}_{n}(v)$ be the $p$-capacities associated with the Galton–Watson tree $\tilde{T}_{n}$ of depth $n$ with reproduction law $\tilde{\mu}$. Since we have truncated the offspring distribution there is a coupling for which $\tilde{T}\subset T$ and since the function $g(x)=\frac{x}{(1+x^{s})^{1/s}}$ is increasing, we obtain that $\tilde{C}_{n}(v)\leqslant C_{n}(v)$ for all $v\in\tilde{T}$. In particular, we only need to obtain a lower bound on $\tilde{C}_{n}$, $\mathbf{E}[\tilde{C}_{n}]$.
@@
~$\tilde{\mathcal{R}}_{p}(\rho\leftrightarrow\partial\tilde{T}_{n})^{s}\leqslant\frac{1}{(\tilde{W}_{n})^{q}}\sum_{k=1}^{n}(\tilde{m}R_{n})^{-ks}\frac{1}{\tilde{m}^{k}}\sum_{|v|=k}\tilde{W}_{n}(v)^{q}\,,$
@@
with $h(x)\sim L(1/x)x^{\alpha-1}$ as $x\downarrow 0$ (recall $s=q-1$).
@@
• We now treat the case $mR_{n}\in[1+\frac{1}{n},2]$. Choose $t_{n}$ such that $h(1/t_{n})=c(mR_{n}-1)$ with $c$ small enough so that $\tilde{\delta}_{n}\leqslant\frac{1}{4}(mR_{n}-1)$. Then, we get that $\tilde{m}R_{n}=mR_{n}(1-\tilde{\delta}_{n})\geqslant 1+\frac{1}{2}(mR_{n}-1)$, using that $(1+x)(1-\frac{1}{4}x)\geqslant 1+\frac{1}{2}x$ for $x\in[0,1]$. We end up with
@@ Appendix A About the iteration of the random cluster model on trees
In this section, we prove Proposition 2.1. For $u\in T_{n}$, let us introduce the events
$A_{u}:=\{\exists\text{ open path from }u\text{ to }\partial T_{n}(u)\text{ inside }T_{n}(u)\}\,.$
@@
Figure 1: Illustration of the identity (A.4) that relates the numbers of clusters in $(\bar{T}_{n}(v))_{v\leftarrow u}$ to the number of clusters in $\bar{T}_{n}(u)$. We have illustrated two cases. On the left, the event $A_{u}$ is not verified i.e. $u\not\leftrightarrow\partial T_{n}(u)$: there are open edges $uv$ (in thick red) only connecting to subtrees $\bar{T}_{n}(v)$ where $A_{v}$ is not verified, and each of these edges reduces the number of clusters by one. On the right, the event $A_{u}$ is verified i.e. $u\leftrightarrow\partial T_{n}(u)$: there are open edges $uv$ (in thick red) connecting to subtrees $\bar{T}_{n}(v)$ where $A_{v}$ is verified and all these open edges reduce the global number of clusters only by one (other open edges connecting to subtrees $\bar{T}_{n}(v)$ where $A_{v}$ is not verified still reduce the number of clusters each by one).
@@
$\begin{split}Z_{u}(A_{u}^{c})&=\mathtt{q}^{2-d_{u}}\mathbb{E}_{p}\Big{[}\mathds{1}_{A_{u}^{c}}\prod_{v\leftarrow u}\mathtt{q}^{K_{v}}\mathtt{q}^{-\mathds{1}_{\{\omega_{uv}=1\}}\mathds{1}_{A_{v}^{c}}}\Big{]}\\ &=\mathtt{q}^{2-d_{u}}\mathbb{E}_{p}\Big{[}\prod_{v\leftarrow u}\mathtt{q}^{K_{v}}\big{(}\mathds{1}_{\{\omega_{uv}=0\}}+\mathtt{q}^{-1}\mathds{1}_{\{\omega_{uv}=1\}}\mathds{1}_{A_{v}^{c}}\big{)}\Big{]}\,,\end{split}$
@@
where we have also written $\gamma=\varphi_{q}(1-\mathtt{p})=\psi_{\mathtt{q}}(\beta)$. To conclude the proof of (2.4), it only remains to observe that $\psi_{\mathtt{q}}^{-1}(x)=-\log\varphi_{\mathtt{q}}(x)$, since $\psi_{\mathtt{q}}(t)=\varphi_{\mathtt{q}}(e^{-t})$ and $\varphi_{\mathtt{q}}^{-1}=\varphi_{\mathtt{q}}$.

### v2

@@ 5.1 Convergence in $L^{1}$
Let us write f(x)=xg(x)f(x)=x-g(x), which is a non-negative function (by concavity of $g$), which verifies $f(x)\leqslant c\min(x^{q},x)$ thanks to (1.9). Then, we can rewrite the above iteration as
@@
which goes to 0 as $k\to\infty$ since $\theta_{q}<q$.
@@
where we have used Lemma 4.3. We now show that the upper bound goes to 0 as $k\to\infty$.
@@
which goes to 0 if $k\to\infty$ sufficiently slowly.
@@
The case $mR_{n}=1$ is trivial since then we have that $a_{k}=k$ and $a_{n}=n$. A similar result holds when $|mR_{n}-1|\leqslant n^{-1}$, since we also have $a_{k}\asymp k$ for all $k\leqslant n$ in that case.
@@
For the last inequality, we have used that $f(x)\leqslant c\,\min(x^{q},x)$, so that for $|v|=j$ we have $\mathbf{E}[f(B_{n}(v))]\leqslant c\mathbf{E}[B_{n}(v)^{q}]\leqslant c^{\prime}(b_{n-j})^{q}$, thanks to Proposition 4.2; recall also that s=q1s=q-1.
@@
[…]
For the last inequality, we have used that $\mathbf{E}[|W-W_{k}|]\leqslant\|W-W_{k}\|_{q}\leqslant cm^{-k(1-\frac{1}{q}\theta_{q})}$, see [27, Prop. 1.3].
Note that the assumption that $\limsup_{n\to\infty}mR_{n}\leqslant 1$ ensures that $k_{n}$ goes to $\infty$ and that $\lim_{n\to\infty}a_{n}=+\infty$, see (1.11). Hence, we can choose $k=\hat{k}_{n}\leqslant k_{n}$ going to $\infty$ sufficiently slowly so that the upper bound in (5.6) goes to zero. This concludes the proof that $(\hat{B}_{n})_{n\geqslant 0}$ converges in $L^{1}$ to $W$. ∎
@@ 5.2 Convergence in $L^{q}$
First, (5.3) and (5.4) show that $(I_{\hat{k}_{n},n})^{q}$ and $(J_{\hat{k}_{n},n})^{q}$ converge to 0 in $L^{1}$; in particular they are uniformly integrable. Second, since $\mu$ admits a finite moment of order $q$, we have that $(W_{k})_{k\geqslant 0}$ converges in $L^{q}$ to $W$, so in particular $(W_{k}^{q})_{k\geqslant 0}$ is uniformly integrable. This concludes the proof that $(\hat{B}_{n}^{q})_{n\geqslant 0}$ is uniformly integrable, hence that $(\hat{B}_{n})_{n\geqslant 0}$ converges in $L^{q}$ to $W$. ∎
@@ 5.3 Almost sure convergence
Note that, setting f(x)=xg(x)f(x)=x-g(x), we have the relation $b_{n}=mR_{n}\big(b_{n-1}-\mathbf{E}[f(B_{n}(v))]\big)$ for $|v|=1$, which can be rewritten as
@@
Now, observe that since $f(x)\sim\kappa_{g}x^{q}$ as $x\downarrow 0$ and that $f(x)\leqslant cx^{q}$, we get by dominated convergence that $\mathbf{E}[f(B_{n}(v))]\sim\kappa_{g}\mathbf{E}[B_{n}(v)^{q}]\sim\kappa_{g}\mathbf{E}[W^{q}](b_{n-1})^{q}$, with $b_{n-1}\downarrow 0$ under the assumption of Proposition 1.9.
In particular, this gives that $c_{n-1}\sim\kappa_{g}\mathbf{E}[W^{q}](b_{n-1})^{s}$, with s=q1s=q-1. Setting $x_{n}=b_{n}^{-s}$, we get that
@@
with $c^{\prime}_{n-1}=(1-c_{n-1})^{-s}-1\sim s\kappa_{g}\mathbf{E}[W^{q}](b_{n-1})^{s}$. Iterating this relation, we get that
@@
with $x_{i}c^{\prime}_{i}\to s\kappa_{g}\mathbf{E}[W^{q}]$ as $i\to\infty$. Hence, we can choose $\ell_{n}$ going to $\infty$ arbitrarily slowly and write that
$\begin{split}x_{n}&=(1+o(1))s\kappa_{g}\mathbf{E}[W^{q}]\sum_{k=1}^{n-\ell_{n}}(mR_{n})^{-ks}+(mR_{n})^{-ns}\sum_{i=0}^{\ell_{n}}(mR_{n})^{is}x_{i}c^{\prime}_{i}\\ &=(1+o(1))s\kappa_{g}\mathbf{E}[W^{q}]a_{n-\ell_{n}}+(mR_{n})^{-ns}\sum_{i=0}^{\ell_{n}}(mR_{n})^{is}x_{i}c^{\prime}_{i}\,.\end{split}$
To conclude, let us notice that, in the case where $\lim_{n\to\infty}mR_{n}=\vartheta=1$, then we can choose $\ell_{n}\to\infty$ so that $a_{n-\ell_{n}}\sim a_{n}$. We also find in that case that the second term is negligible compared to $a_{n}$ provided that $\ell_{n}$ grows sufficiently slowly — this is clear if $mR_{n}\geqslant 1-\frac{c}{n}$ and follows from the fact that $a_{n}\sim(1-(mR_{n})^{s})^{-1}(mR_{n})^{-(n+1)s}$ if $n(mR_{n}-1)\to$-$\infty$. This shows that $x_{n}\sim s\kappa_{g}\mathbf{E}[W^{q}]a_{n}$ as $n\to\infty$ when $\vartheta=1$.
@@ 6 Estimates on moments of $C_{n}$ in the case $\mathbf{E}[Z^{q}]=+\infty$
In this section, we prove Theorem 1.10. Since we can bound $B_{n}$ with $C_{n}:=\mathcal{C}_{p}(\rho\leftrightarrow\partial T_{n})$, see Proposition 1.4, we focus on estimates on $C_{n}$, $\mathbf{E}[C_{n}]$. Let us define, as in Section 4.1, $\phi(u)=R_{u}\mathcal{C}_{p}(u\leftrightarrow\partial T_{n}(u))$ and recall that we have the following recursion (4.1):
@@ 6.1 Upper bound on $\mathbf{E}[C_{n}]$
Let us assume that (1.13) holds and obtain the upper bound in Theorem 1.10. As in Section 4.2, let $\varphi_{k}=\mathbf{E}[\phi(u)]$ for $|u|=k$, so in particular $\varphi_{n}=1$ and we need to estimate $\varphi_{0}$. Using the recursion (6.1) we can write, again with the notation f(x)=xg(x){f(x)=x-g(x)},
[…]
@@
Notice that the recursion (6.4) admits a non-zero fixed point $u_{*}$ if one has $mR_{n}>1$, and that it verifies $h(u_{*})=\frac{mR_{n}-1}{mR_{n}}$. If on the other hand $mR_{n}\leqslant 1$, the only fixed point is $u_{*}=0$.
• Let us start with the case where $mR_{n}\in[1+\frac{1}{n},K]$, and let $\delta_{n}:=\frac{mR_{n}-1}{mR_{n}}$. We let $\upsilon_{n}$ be the fixed point of the equation (6.4), i.e. such that $h(\upsilon_{n})=\delta_{n}$, and notice that it verifies $\upsilon_{n}=h^{-1}(\delta_{n})\leqslant c\gamma_{n}$, where $\gamma_{n}$ is defined in (1.15). Let us assume that $\upsilon_{n}\leqslant u_{0}$, otherwise we have $u_{k}\leqslant\upsilon_{n}$ for all $k$ and in particular $u_{n}\leqslant\upsilon_{n}\leqslant c\gamma_{n}$.
@@
Now, applying this inequality with $k=k_{n}-1$ and recalling the definition of $k_{n}$, we get that $k_{n}-1\leqslant c^{\prime\prime}C^{-1}/\delta_{n}$. Since $\delta_{n}=(mR_{n}-1)/mR_{n}\geqslant 1/Kn$ for $mR_{n}\in[1+\frac{1}{n},K]$, we get that $k_{n}-1\leqslant Kc^{\prime\prime}C^{-1}n$, which is smaller than $n$ provided that $C$ had been fixed large enough.
@@ 6.2 Lower bound on $C_{n}$, $\mathbf{E}[C_{n}]$
To obtain a lower bound on $C_{n}$, $\mathbf{E}[C_{n}]$, let us assume that (1.14) holds. We let $(t_{n})_{n\geqslant 1}$ be a truncation sequence (to be optimized later on), and as in 3.2, we consider a truncated branching process $\tilde{T}$ with offspring distribution $\tilde{Z}:=(Z\wedge t_{n})\sim\tilde{\mu}$. Let also $\tilde{C}_{n}(v)$ be the $p$-conductances associated with the Galton–Watson tree $\tilde{T}_{n}$ of depth $n$ with reproduction law $\tilde{\mu}$. Since we have truncated the offspring distribution there is a coupling for which $\tilde{T}\subset T$ and since the function $g(x)=\frac{x}{(1+x^{s})^{1/s}}$ is increasing, we obtain that $\tilde{C}_{n}(v)\leqslant C_{n}(v)$ for all $v\in\tilde{T}$. In particular, we only need to obtain a lower bound on $\tilde{C}_{n}$, $\mathbf{E}[\tilde{C}_{n}]$.
@@
$\tilde{\mathcal{R}}_{p}(\rho\leftrightarrow\partial\tilde{T}_{n})^{s}\leqslant\frac{1}{(\tilde{W}_{n})^{q}}\sum_{k=1}^{n}(\tilde{m}R_{n})^{-ks}\frac{1}{\tilde{m}^{k}}\sum_{|v|=k}\tilde{W}_{n}(v)^{q}\,$,
@@
with $h(x)\sim L(1/x)x^{\alpha-1}$ as $x\downarrow 0$ (recall s=q1s=q-1).
@@
• We now treat the case $mR_{n}\in[1+\frac{1}{n},K]$. Choose $t_{n}$ such that $h(1/t_{n})=c(mR_{n}-1)$ with $c$ small enough so that $\tilde{\delta}_{n}\leqslant\frac{1}{4}(mR_{n}-1)$. Then, we get that $\tilde{m}R_{n}=mR_{n}(1-\tilde{\delta}_{n})\geqslant 1+\frac{1}{2}(mR_{n}-1)$, using that $(1+x)(1-\frac{1}{4}x)\geqslant 1+\frac{1}{2}x$ for $x\in[0,1]$. We end up with
@@ Appendix A About the iteration of the random cluster model on trees
In this section, we prove Proposition 2.1. For $u\in T_{n}$, let us introduce the event
[…]
@@
Figure 1: Illustration of the identity (A.4) that relates the numbers of clusters in $(\bar{T}_{n}(v))_{v\leftarrow u}$ to the number of clusters in $\bar{T}_{n}(u)$. We have illustrated two cases. On the left, the event $A_{u}$ is not verified i.e. $u\not\leftrightarrow\partial T_{n}(u)$: there are open edges uvuv (in thick red) only connecting to subtrees $\bar{T}_{n}(v)$ where $A_{v}$ is not verified, and each of these edges reduces the number of clusters by one. On the right, the event $A_{u}$ is verified i.e. $u\leftrightarrow\partial T_{n}(u)$: there are open edges uvuv (in thick red) connecting to subtrees $\bar{T}_{n}(v)$ where $A_{v}$ is verified and all these open edges reduce the global number of clusters only by one (other open edges connecting to subtrees $\bar{T}_{n}(v)$ where $A_{v}$ is not verified still reduce the number of clusters each by one).
@@
[…]
@@
where we have also written $\gamma=\varphi_{\mathtt{q}}(1-\mathtt{p})=\psi_{\mathtt{q}}(\beta)$. To conclude the proof of (2.4), it only remains to observe that $\psi_{\mathtt{q}}^{-1}(x)=-\log\varphi_{\mathtt{q}}(x)$, since $\psi_{\mathtt{q}}(t)=\varphi_{\mathtt{q}}(e^{-t})$ and $\varphi_{\mathtt{q}}^{-1}=\varphi_{\mathtt{q}}$.